OpenStudy (anonymous):

A rectangular garden measures 34 by 45 . Surrounding (and bordering) the garden is a path 2ft wide. Find the area of this path. Be sure to include the correct unit in your answer.

OpenStudy (btaylor):

 The Area of the Path is (38 x 49) - (34 x 45) = 1862 - 1530 = 332